OpenStudy (anonymous):

3[-x + (2x+1)] =x - 1

OpenStudy (anonymous):

-3X+6X+3=X-1 THEN we have: -3x+6x-x=-1-3 then we have: 2x=-4 then x=-4/2 then x=-2

OpenStudy (anonymous):

@og.dac